OpenStudy (anonymous):

A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red?

OpenStudy (zarkon):

since this is sampling with replacement these events are independent.

Parth (parthkohli):

\[\Large {1 \over 100} \] This is the probability as the denominator is squared. \[\Large {1 \over 10*10} = {1 \over 100}\]

OpenStudy (anonymous):

thx guys :)

OpenStudy (zarkon):

that is not correct

Parth (parthkohli):

Oops yes, I'm wrong

OpenStudy (anonymous):

:) THERE R THE POSSIBLE ANSWERS A . 3/5 B . 7/20 C . 3/25 D . 7/100

Parth (parthkohli):

C is the answer. I'm completely sure.

OpenStudy (anonymous):

thx soooooo much :)

OpenStudy (zarkon):

\[\frac{3}{10}\frac{4}{10}=\frac{12}{100}=\frac{3}{25}\]

Parth (parthkohli):

\[\Large {12 \over 100} = {3 \over 25} = C\]

OpenStudy (anonymous):

got it Thank you...